Suppose Betty's demand for physician visits is Quantity = 10-(0.2*Price).
a. Draw Betty's demand curve.
b. What is the quantity demanded at a price of $10 per visit? $25?
c. At what price will she buy four visits? Eight visits?
d. If the government agrees to pay half of her health care bills, what would her quantity demanded be at a price of $10 per visit? $25? Draw the new,subsidized demand curve.
Answer #1

A.

55 -50 -45 -40 35 30 25 Bettys-Demand-curve- -15 10 -5 2. 10 0. 4 9. 8. Quantity -5 Price 20

B. When Price= $10 per visit

Quantity= 10-0.2*10= 8

When Price= $25 per visit

Quantity= 10-0.2*25= 5

C. Quantity= 4 visits

4= 10-0.2* Price

Price= $30

Quantity= 8 visits

8= 10-0.2*Price

Price= $10

D. New subsidized demand curve equation:

Quantity= 10-0.2*0.5*Price

Price= $10

Quantity= 10-0.2*0.5*10= 9

Price= $25

Quantity= 10-0.2*0.5*25= 7.5

120- 110 -100 06- -80 70 60 50 -40- -30 New subsidized-Demand 20 10 14 16 4. 10 12 -2 2 8. Quantity -10- Price 6.

If it helps kindly update.